# Predicting Open-Hole Laminates Failure Using Support Vector Machines With Classical and Quantum Kernels

### v1

@@ Abstract
Modeling open hole failure of composites is a complex task, consisting in a highly nonlinear response with interacting failure modes. Numerical modeling of this phenomenon has traditionally been based on the finite element method, but requires to tradeoff between high fidelity and computational cost. To mitigate this shortcoming, recent work has leveraged machine learning to predict the strength of open hole composite specimens. Here, we also propose using data-based models but to tackle open hole composite failure from a classification point of view. More specifically, we show how to train surrogate models to learn the ultimate failure envelope of an open hole composite plate under in-plane loading. To achieve this, we solve the classification problem via support vector machine (SVM) and test different classifiers by changing the SVM kernel function. The flexibility of kernel-based SVM also allows us to integrate the recently developed quantum kernels in our algorithm and compare them with the standard radial basis function (RBF) kernel. Finally, thanks to kernel-target alignment optimization, we tune the free parameters of all kernels to best separate safe and failure-inducing loading states. The results show classification accuracies higher than 90% for RBF, especially after alignment, followed closely by the quantum kernel classifiers.
@@ 1 Introduction
The rest of the paper is structured as follows. Section 2 describes the machine learning problem, by defining the input, the data sampling strategy and the labeling criterion. Section 3 briefly introduces the SVM dual problem, the RBF kernels and the quantum kernels. More details about these methods are available in the appendices following the main body of the manuscript. Finally, Section 4 presents the classification results for all kernels and Section 5 outlines conclusions and future work.
@@ 3 Methodology
Figure 4: Quantum embeddings. (a) Generic quantum embedding made of one- and two-qubit gates. The width of the embedding is the number of qubits, while the depth is the number of layers, which is a minimal block of gates. (b) IQP quantum embedding layer, which is parameter free, but encodes products of features as $ZZ$ interactions. (c) HE2 quantum embedding layer, parametrized by $\left[\theta_{0},\theta_{1},\dots,\theta_{P-1}\right]$
@@ 4 Results
We tested our machine learning models on a dataset of 1960 labelled strain vectors $\bm{\varepsilon}^{(m)}$, which we obtained by uniformly sampling the homogeneous strain/stress pairs from the FE simulations of the open-hole composite specimen. The input homogeneous strains in both normal and shear directions were varied between ${1}\mathrm{e}{-4}$ and ${1}\mathrm{e}{4}$ microstrains and a stiffness degradation threshold of 0.9 was used to discriminate non-failed and failed loading states.
@@
The same analysis was performed for all the quantum kernels considered, where we wanted to take into account the effect on accuracy of different embeddings and of maximizing the kernel-target alignment. The results are reported in Figure 8, which shows accuracies roughly between 67% and 87% for all embeddings with different values of $C$. Except for IQP case, increasing $C$ leads to higher accuracies, hinting to the the need of a tight bound when mapping with these embeddings, similar to the RBF kernel. Unfortunately, for $C\geq{1}\mathrm{e}{4}$, the optimization of the dual SVM failed to converge for the quantum kernels, likely due to numerical ill-conditioning, presumably preventing from reaching higher accuracies. In fact, we observe that the accuracies of both untrained and trained HE2 kernels monotonically increase with $C$. For the trained HE2 case, this is true regardless of the number of qubits and depth. Furthermore, Figure 8 also shows that increasing the embedding resources, especially the number of qubits, pays off more when also optimizing the KTA.
Figure 9: Test data classification accuracy of classical and quantum kernels for increasing training set size. The RBF kernel-SVMs were trained with both $C={1}\mathrm{e}{7}$ and $C={1}\mathrm{e}{4}$, as the latter is the highest value of $C$ for which the quantum-kernel SVMs could still be fitted. Quantum-kernel classifiers were trained instead with the embedding arhitecture and $C$ value ensuring highest accuracy during grid-search cross validation
Classical and quantum kernels are finally compared in Figure 9, which shows how 5 different models classify a test set of strain loading data when fitted on progressively larger training sets. A similar comparison on additional classification metrics can be found in Appendix D. The RBF kernel achieves 80% accuracy with just 10% the total training set size, and with $C={1}\mathrm{e}{7}$ it reaches over 90% with just half the training points. In comparison, all quantum kernel classifiers are at least 5% less accurate than the best RBF-kernel SVM. However, especially for HE2 embeddings, the scores are similar to the $C={1}\mathrm{e}{4}$ RBF case, suggesting that RBF and HE2 kernels separate the non-failed and failed classes to a similar extent. Changing the embedding from HE2 to IQP, there is a drop in accuracy for small training set sizes, while the performance is similar when more than half the training set is used. On the other hand, the effect of training the kernel is less visible at this stage, reflecting the fact that the accuracies obtained during grid-search cross validation are alike for untrained and trained HE2.
@@ Appendix D Classical and quantum SVM comparison on different classification metrics
Table 1: Comparison of five different classification scores between the RBF-SVM and the trained HE2W6D3 quantum kernel SVM. The classifiers were trained with increasing fractions of the training dataset. Notice that the RBF-SVM problem used $C={1}\mathrm{e}{7}$, while the HE2W6D3 quantum kernel SVM used $C={1}\mathrm{e}{4}$, which is the highest $C$ values before the occurrence of convergence issues.

### v2

@@ Abstract
Modeling open hole failure of composites is a complex task, consisting in a highly nonlinear response with interacting failure modes. Numerical modeling of this pheonomenon has traditionally been based on the finite element method, but requires to tradeoff between high fidelity and computational cost. To mitigate this shortcoming, recent work has leveraged machine learning to predict the strength of open hole composite specimens. Here, we also propose using data-based models but to tackle open hole composite failure from a classification point of view. More specifically, we show how to train surrogate models to learn the ultimate failure envelope of an open hole composite plate under in-plane loading. To achieve this, we solve the classification problem via support vector machine (SVM) and test different classifiers by changing the SVM kernel function. The flexibility of kernel-based SVM also allows us to integrate the recently developed quantum kernels in our algorithm and compare them with the standard radial basis function (RBF) kernel. Finally, thanks to kernel-target alignment optimization, we tune the free parameters of all kernels to best separate safe and failure-inducing loading states. The results show classification accuracies higher than 90% for RBF, especially after alignment, followed closely by the quantum kernel classifiers.
@@ 1 Introduction
The rest of the paper is structured as follows. Section 2 describes the machine learning problem, by defining the input, the data sampling strategy and the labeling criterion. Section 3 briefly introduces the SVM dual problem, the Radial Basis Functions (RBF) kernels and the quantum kernels. More details about these methods are available in the appendices following the main body of the manuscript. Finally, Section 4 presents the classification results for all kernels and Section 5 outlines conclusions and future work.
@@ 3 Methodology
Figure 4: Quantum embeddings. (a) Generic quantum embedding made of one- and two-qubit gates. The width of the embedding is the number of qubits, while the depth is the number of layers, which is a minimal block of gates. (b) IQP quantum embedding layer, which is parameter free, but encodes products of features as $ZZ$ interactions. (c) HE2 quantum embedding layer, parametrized by $\left[\theta_{0},\theta_{1},\dots,\theta_{P-1}\right]$.
@@ 4 Results
We tested our machine learning models on a dataset of 1960 labelled strain vectors $\bm{\varepsilon}^{(m)}$, which we obtained by uniformly sampling the homogeneous strain/stress pairs from the FE simulations of the open-hole composite specimen. The input homogeneous strains in both normal and shear directions were varied between $10^{-4}$ and $10^{4}$ microstrains and a stiffness degradation threshold of 0.9 was used to discriminate non-failed and failed loading states.
@@
The same analysis was performed for all the quantum kernels considered, where we wanted to take into account the effect on accuracy of different embeddings and of maximizing the kernel-target alignment. The results are reported in Figure 8, which shows accuracies roughly between 67% and 87% for all embeddings with different values of $C$. Except for IQP case, increasing $C$ leads to higher accuracies, hinting to the the need of a tight bound when mapping with these embeddings, similar to the RBF kernel. Unfortunately, for high values of $C$, the optimization of the dual SVM failed to converge for some of the quantum kernels, likely due to numerical ill-conditioning. This presumably prevented the quantum kernel classifiers from even better separating failed instances, as suggested by the monotonic increasing test accuracies with $C$, at least for the HE2 kernels. Furthermore, Figure 8 shows that the scores improve when more embedding resources (number of qubits and layers) are added, especially in the case of KTA-optimized HE2 kernel.
Figure 9: Test data classification accuracy of classical and quantum kernels for increasing training set size. The RBF kernel-SVMs were trained with both $C=10^{7}$ and $C=10^{4}$, as the latter is the highest value of $C$ for which the quantum-kernel SVMs could still be fitted. Quantum-kernel classifiers were trained instead with the embedding arhitecture and $C$ value ensuring highest accuracy during grid-search cross validation.
Classical and quantum kernels are finally compared in Figure 9, which shows how 5 different models classify a test set of strain loading data when fitted on progressively larger training sets. A similar comparison on additional classification metrics can be found in Appendix D. The RBF kernel achieves 80% accuracy with just 10% the total training set size, and with $C=10^{7}$ it reaches over 90% with just half the training points. In comparison, all quantum kernel classifiers are at least 5% less accurate than the best RBF-kernel SVM. However, especially for HE2 embeddings, the scores are similar to the $C=10^{4}$ RBF case, suggesting that RBF and HE2 kernels separate the non-failed and failed classes to a similar extent. Changing the embedding from HE2 to IQP, there is a drop in accuracy for small training set sizes, while the performance is similar when more than half the training set is used. On the other hand, the effect of training the kernel is less visible at this stage, reflecting the fact that the accuracies obtained during grid-search cross validation are alike for untrained and trained HE2.
@@ Appendix D Classical and quantum SVM comparison on different classification metrics
Table 1: Comparison of five different classification scores between the RBF-SVM and the trained HE2W6D3 quantum kernel SVM. The classifiers were trained with increasing fractions of the training dataset. Notice that the RBF-SVM problem used $C=10^{7}$, while the HE2W6D3 quantum kernel SVM used $C=10^{4}$, which is the highest $C$ values before the occurrence of convergence issues.